OpenStudy (celinegirl):

Hi please check my answers.. I want to make sure Part A and B are correct. Can someone also help with part C?

OpenStudy (celinegirl):

Here is the problem: Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

your answer is correct

OpenStudy (celinegirl):

@mya1234567890 I havent even posted it yet

OpenStudy (celinegirl):

Here is my answer: Part A: x = days and y = total rent Two points: (5,465) and (7,625) Find the slope using those two points. (625 - 465) / (7 - 5) = 160/2 = 80 Then use m = 80 and one of the points to write the equation of the line in point-slope form: y - y(1) = m(x - x(1) y - 465 = 80(x - 5) convert: y = 80(x - 5) = 465 y = 80x - 400 = 456 y = 80x = 65 7 - 80x = 65 Part B: f(x) 80x + 65

OpenStudy (anonymous):

okay i belive that

OpenStudy (bruhobamacare):

A and B are correct

OpenStudy (celinegirl):

Thank you. Does anyone know part C?

OpenStudy (bruhobamacare):

just use x-values to represent the days and the output will represent the cost. So for example if you plug in 1 in the formula, you would get 80 + 65 = 145 $ which is the cost of one night. then 80 + 130 = 210, which is the cost of the second night and so on. After plugging in all the points you should notice that this is a linear function.. if this makes any sense :)

OpenStudy (celinegirl):

Hold on I got something different for part A: 80x - y = -65

OpenStudy (celinegirl):

Part B is the same. Part C: f(x) = 80x =65 solve for f(6) First substitute x for 6 f(6) = 80(6) + 65 next, solve 80(6) f(6) = 480 = 65 then add 64 f(6) = 545 (6, 545) Solve for f(4) First, subsistute x for 4 f(4) = 80(4) = 65 Next, solve 80(4) f(4) = 320 = 65 then add 64 f(4) = 385 (4, 385)

OpenStudy (celinegirl):

Would that be correct @BruhObamaCare

OpenStudy (bruhobamacare):

yes :)

OpenStudy (celinegirl):

@BruhObamaCare thank you so much!

OpenStudy (bruhobamacare):

anytime :)